### v4

# Exploring the Low-Pass Filtering Behavior in Image Super-Resolution

Abstract
Deep neural networks for image super-resolution (ISR) have shown significant advantages over traditional approaches like the interpolation. However, they are often criticized as ‘black boxes’ compared to traditional approaches with solid mathematical foundations. In this paper, we attempt to interpret the behavior of deep neural networks in ISR using theories from the field of signal processing. First, we report an intriguing phenomenon, referred to as ‘the sinc phenomenon.’ It occurs when an impulse input is fed to a neural network. Then, building on this observation, we propose a method named Hybrid Response Analysis (HyRA) to analyze the behavior of neural networks in ISR tasks. Specifically, HyRA decomposes a neural network into a parallel connection of a linear system and a non-linear system and demonstrates that the linear system functions as a low-pass filter while the non-linear system injects high-frequency information. Finally, to quantify the injected high-frequency information, we introduce a metric for image-to-image tasks called Frequency Spectrum Distribution Similarity (FSDS). FSDS reflects the distribution similarity of different frequency components and can capture nuances that traditional metrics may overlook. Code, videos and raw experimental results for this paper can be found in: https://github.com/RisingEntropy/LPFInISR.
Machine Learning, ICML
Please refer to Tab. 2 for notation conventions.
1 Introduction
Figure 1: $I$ is an image in which only the central pixel is 1 and the other pixels are 0. What would the result look like if image I is super-resolved using a neural network, A, B, C, or D? Surprisingly, the answer is A. We name this phenomenon as the sinc phenomenon. In this paper, we give a possible explanation for this phenomenon.
The goal of image super-resolution (ISR) is to reconstruct low-resolution (LR) images into high-resolution (HR) images through various techniques. In recent years, with advances in deep learning, growing ISR methods using neural networks are proposed, bringing the development of ISR into a new level. While impressive results persistently arise, the mechanism under ISR networks remain largely unexplored, leading to criticism that they are considered black boxes. In comparison, traditional methods, such as interpolation or filtering, have strong interpretability. Despite the principles of traditional methods and neural networks are different, we can still attempt to explain the behavior of ISR networks using theories from traditional methods. In this paper, following this line of thought, we successfully utilize theories from the field of signal processing techniques to explain the performance of neural networks in the ISR task.
The target of the ISR task is to upsample a two-dimensional signal. In traditional signal processing theory [17, 18], a feasible method for upsample involves restoring a discrete low-sampling-rate signal to a continuous signal using a low-pass filter, and then sampling the continuous signal at a higher rate to obtain a high-sampling-rate signal. An intriguing aspect of this process is that when we try to upsample a Dirac $\delta$ signal, we will finally get a sinc signal since the sinc signal is the time-domain waveform of a low-pass filter, (for details about this, please refer to Sec. 3 ). Given this, we can conjecture: if neural networks exhibit similar behavior, then when attempt to super-resolve a Dirac $\delta$ signal, the resultant outcome would also be a sinc signal. As shown in Fig. 1, we indeed observe this phenomenon, and we name it as ‘the sinc phenomenon’. This phenomenon establishes a connection between traditional signal processing theory and the interpretability of neural networks, thus helping us form a deeper understanding of ISR networks.
[…]
Given that the non-linear component is injecting high-frequency information, there is a pressing need for a metric to quantitatively describe the extent of the injected high frequencies. Previous metrics, like PSNR, SSIM [26] and LPIPS [36], have not approached the evaluation of images from a frequency perspective. Therefore, we propose the frequency spectrum distribution similarity (FSDS), a metric that evaluates image quality based on the power distribution in the frequency spectrum.
In summary, our contribution can be concluded as:
  • We report an intriguing phenomenon: the impulse responses of image super-resolution (ISR) networks are sinc functions, representing the temporal waveform of a low-pass filter. We name it the ’sinc phenomenon’. This observation helps to establish a connection between signal processing theory and neural networks. Moreover, we find that for a network, the more similar the impulse response is to the sinc function, the better performance it produces.
  • In order to further explain the performance of neural networks in the ISR task through this phenomenon, we introduce HyRA. HyRA considers the neural network as a parallel combination of a linear system and a non-linear system with a zero impulse response. It points out that the linear system operates as a low-pass filter, while the non-linear system injects high-frequency information.
  • To quantitatively describe the injection of high frequencies, we introduce the FSDS metric. FSDS measures image quality using frequency spectrum produced by FFT and can reflect high-frequency distortions that previous metrics fail to capture.
2 Related works
2.1 Super Resolution Using Neural Networks
Recent review articles in ISR include fixed-scale super-resolution [33] and arbitrary-scale super-resolution review [14]. There are various architectures of mainstream ISR backbone networks, including CNN-style backbones [13, 38, 9, 37, 2], transformer-style backbones [12, 25] and GAN-style backbone networks [24], etc. Based on these backbones, researchers have proposed quantitative modules with various functions. For example, ArbSR [23] can expand a fixed-scale super-resolution network to an arbitrary-scale ISR network, LTE [10] can enhance local textures, etc. What worth mentioning is that LIIF [5] introduces implicit neural representation into ISR for the first time, bringing a new approach for ISR. This paper mainly focuses on approaches that utilize CNN-style or transformer-style backbones. Except for network architectures, numerous datasets have been proposed to facilitate further research. Commonly used datasets for ISR includes Set5 [3], Urban100 [8], Flickr2K [34], SCI1K [32], DIV2K [1], etc. We evaluate the effectiveness of our proposed FSDS metric on DIV2K dataset. The large size of the DIV2K dataset contributes to increased reliability in our conclusions.
2.2 Explaining the Behavior of Neural Networks
Despite neural networks are often criticized as ‘black boxes,’ predecessors have made remarkable efforts to mitigate this situation. Various previous researches have proposed plenty of methods to analyze the behavior of neural networks. Since Sundararajan et al. [21] introduce the integrated gradients (IG) for attribution in classification tasks, numerous researchers have expanded this method to various domains, broadening the scope of attribution beyond classification tasks. Based on IG, Gu & Dong [6] propose LAM to analyze the impact of the local patch on the entire ISR outcome. However, such a method requires manually determined hyper-parameters and baselines, thus introducing subjectivity. Several notable analysis methods utilizing the Fourier transform have been explored in the literature [31, 29, 30, 39]. Notably, Xu et al. [31] propose the Frequency-Principle, claiming its relevance to both convolutional neural networks (CNNs) and fully-connected deep neural networks. According to their proposition, these networks inherently adhere to the Frequency-Principle, wherein training data is systematically acquired in a sequential manner, progressing from low to high frequency. Unlike previous approaches these approaches, HyRA distinguishes itself by employing impulse response to probe the potential mechanisms of deep neural networks in the context of the ISR task.
3 Preliminaries
Section B.1-Section B.3 provide a brief overview of signal processing concepts for readers who are not familiar with it. Section B.1 introduces the concepts of signals and systems, along with the computation of responses in Linear Time-Invariant (LTI) systems. Section B.2 covers the processes of signal sampling and reconstruction. Section B.3 delves into the phenomenon of spectrum aliasing, a factor contributing to the ill-posed nature of the ISR task. And we will introduce applying low-pass filter for ISR here.
We can employ signal recovery methods to achieve image super-resolution (ISR). Initially, we conceptualize an image as a series of impulse trains in a two-dimensional continuous space, with varying densities representing different resolutions. Then, for the low-resolution image, we begin by implementing low-pass filtering, following the procedure outlined in Section B.2, to obtain the continuous image $I^{\mathrm{cont}}$, This process can be mathematically described as:
$I^{\mathrm{cont}}_{x,y}=sinc^{\mathrm{\omega}}_{x,y}*I^{\mathrm{LR}}_{x,y},$ (1)
where * denotes convolution, $I^{LR}_{x,y}$ is the low resolution image with variant $x,y$ and $I^{\mathrm{cont}}_{x,y}$ is the continuous signal. $sinc^{\omega}_{x,y}$ is a two-dimensional sinc function with parameter $\omega$ [^2] Please refer to Tab. 3 for Fourier transform pairs, whose frequency spectrum is an ideal low-pass filter with a passband of $0\sim\omega$. Subsequently, we sample the ‘conceptually continuous signal’ at an elevated sampling rate to acquire a more densely populated two-dimensional sequence of impulse trains, i.e., an image with higher resolution denoted as $I^{SR}$:
$I^{SR}_{x,y}=I^{\mathrm{cont}}_{x,y}\cdot s^{\Delta X,\Delta Y}_{x,y}.$ (2)
In the equation, $s^{\Delta X,\Delta Y}_{x,y}$ denotes the two-dimensional impulse trains with intervals of $\Delta X$ in $x$ axis and $\Delta Y$ in $y$ axis.
Figure 2: Various interpolation kernels for ISR. They can all be seen as an approximation of sinc function.
In fact, commonly used interpolation kernels for ISR, such as nearest-neighbor interpolation, linear interpolation, cubic interpolation, etc., can be seen as approximations of the sinc function considering a balance between computational complexity and effectiveness, as illustrated in Fig. 2. Taking into account the similarity of these interpolation kernels, in this paper, we collectively refer to these parameter-free methods as low-pass filter-based super-resolution methods.
4 Method
4.1 Hybrid Response Analysis (HyRA)
Figure 3: Conceptual diagram of HyRA’s core idea.
In this section, we describe the proposed Hybrid Response Analysis (HyRA), which treats the neural network as a combination of a linear system and a non-linear system. Through the impulse response, we can calculate a linear time invariant (LTI) system’s output from any input using the convolution operation (see Section B.1). However, since neural networks are nonlinear systems, we cannot apply convolution to analyze them. To further explore the network features, we need to split it into a linear system and a non-linear system, i.e., HyRA. The core concept HyRA is illustrated in Fig. 3. We denote an ISR network as $N(I)$, where $I$ is the input image. $N(I)$ is a non-linear system that can be expressed as the sum of a linear system and a non-linear system:
$N(I)=H(I)+G(I).$ (3)
In the equation, $H(I)$ represents a linear system, and $G(I)$ represents a non-linear system. Without constraints, such a representation is meaningless because $H(I)$ can be arbitrarily chosen, leading to an infinite variety of representations with the same form but different meanings. To give meaning to this representation, we introduce a constraint: the impulse response of $G(I)$ is zero. With this constraint, both $H(I)$ and $G(I)$ can be uniquely determined. Lemma 4.1 demonstrates that under this constraint, $N(I)$ can still be expressed in the form of Eq. 3. This straightforward method is the essence of HyRA.
For the ISR task, there is a distinctive property known as a ‘spatially invariant system’ [15] associated with it. Consider the definition of time-invariant systems as mentioned in Section B.1, we can naturally extend the concept of in-variance from one-dimensional to two-dimensional space and the definition of spatially invariant systems is: when the input is $I_{x,y}$, the output is $G(I_{x,y})=O(x,y)$; when the input becomes $I^{\prime}=I_{x-x_{0},y-y_{0}}$, the output should be $G(I^{\prime})=O(x-x_{0},y-y_{0})$. For convolution based architectures, we can easily prove its spatial invariance (see the proof below). For transformer-based architectures, we can still use experiments to prove the spatial invariance (see Fig. 15).
Proof.
A convolution operation can be defined as:
$Conv_{i,j}=\sum_{p,q}I_{i-p,j-q}K_{p,q}.$
Then, the shifting operation can be defined as:
$\operatorname{Sh}(i,j)\to(i+k,j+l).$
Combine these two, we then have:
$Conv_{\operatorname{Sh}(i,j)}$ $=Conv_{i+k,j+l}$
$=\sum_{p,q}I_{i+k-p,j+l-q}K_{p,q}$
$=\operatorname{Sh}(\sum I_{i-p,j-q}K_{p,q})$
$=\operatorname{Sh}(Conv_{i,j}).$
This is the invariance of a single convolution layer, and still holds for more layers. ∎
According to HyRA, when we input a Dirac $\delta$ signal to the neural network, we can get the impulse response of the linear system (please recall Section B.1), denoted as $H(\delta)$. For any input $I$, the response of the linear space invariant system can be obtained by convolving the input with the obtained impulse response, which can be expressed as:
$H(I)=I*H(\delta),$ (4)
where * means the convolution operation. Although the response of the non-linear component cannot be directly computed, if we obtain the final output of the neural network, the non-linear part can be deduced by subtracting the response of the linear component from the final output, namely the non-linear response can be computed as:
$G(I)$ $=N(I)-H(I)$ (5)
$=N(I)-I*H(\delta).$
Lemma 4.1.
A neural network $N(I)$ can be expressed as a combination of a linear system $H(I)$ and a non-linear system with an impulse response of zero, i.e., $N(I)=H(I)+G(I)$, where $G(\delta)=0$. Here, $\delta$ represents the Dirac delta function.
Proof.
1) When $G(\delta)=0$, the conclusion holds.
2) When $G(\delta)\neq 0$, Let $H_{1}(I)=H(I)+G(\delta)*I$ and $G_{1}(I)=G(I)-G(\delta)*I$. In this case, $H_{1}(I)$ remains a linear system and $G^{\prime}(I)$ remains a non-linear system. The equation $N(I)=H_{1}(I)+G_{1}(I)$ holds, and it satisfies $G_{1}(\delta)=0$. ∎
4.1.1 $H(I)$ is a low-pass filter
Figure 4: Top row: a super-resolved image by [13] can be viewed as the summation of a linear response obtained by convolving impulse response with the input and the non-linear response gained by subtracting linear-part from the ISR result. Second row: the corresponding frequency spectrum amplitude of the top row. Third row: the corresponding frequency spectrum phase of the top row. The phase compensation indicates that the non-linear part is compensating distortion.
In Sec. 3, we mention that a simple low-pass filter achieves ISR functionality. Do neural networks possess low-pass filters internally? If this hypothesis is valid, according to the principle of HyRA, when we input a Dirac $\delta$ signal into the neural network $N(I)$, the output should be the impulse response of the low-pass filter, i.e., the sinc function (please recall Section B.1 and Tab. 3). In the experiment section (Section 5.2), we conduct tests on three mainstream ISR backbones and some derived methods. We find that their impulse responses are sinc functions [^3] Strictly speaking, it is a windowed sinc function. Regarding the windowing operation, please refer to Appx. E. Now, with both the impulse response and spatial invariance property, we can compute the response of the linear system $H(I)$ to any input through convolution:
$H(I)_{x,y}$ $=I_{x,y}*H(\delta)$ (6)
$=\iint_{(\tau,u)\in\mathbb{R}^{2}}I_{\tau,u}H(\delta)_{x-\tau,y-u}\mathrm{d}\tau\mathrm{d}u.$
In a practical scenario, when dealing with a two-dimensional impulse array represented by $I$, the integration process can be effectively substituted with summation, incorporating appropriate padding. Despite the convolution operator in PyTorch [19] being inherently a correlation operator, the symmetric nature of the sinc function allows for its seamless utilization within such an operator. We present a toy example in Fig. 4 in which we compute the response of the linear component of the EDSR network [13] during ISR. Observing the experimental results, we notice that the linear function $H(I)$ essentially achieves super-resolution, but there are some issues: edge blurring and the presence of grid-like distortions.
The edge is blurred because the low-pass filter removes some high-frequency details. In the frequency spectrum, it is manifested as a relatively small range of diffusion of the central bright spot towards the surroundings. This implies that the image has more low-frequency components and fewer high-frequency components. Such an outcome is the inevitable consequence of applying the low-pass filter.
When computing the response of the linear system, we first perform zero-interpolation on the low-resolution image to achieve the target spatial size. This operation leads to periodic extension in the frequency spectrum [^4] Please refer to Appx. F in the Appendix for details about the periodic extension.. Since this low-pass filter is not a complete ideal filter, but an ideal filter truncated by a certain window function, its filtering performance is weakened by the window function. The weakened filter cannot completely eliminate the extended spectrum, meaning the attenuation in the stopband is insufficient, as referred to in signal processing, thus causing such gird-like distortions.
In summary, the linear system $H(I)$ (the low-pass filter approximated by the neural network) can achieve super-resolution functionality, but it is not perfect. On one hand, the low-pass filter determines that the image is blurred, lacking high frequencies. On the other hand, the filter is windowed, leading to a weakened filtering performance and resulting in grid-like distortions. These issues will be compensated for by the nonlinear system $G(I)$.
4.1.2 $G(I)$ Injects high-frequency information
Figure 5: An illustration of how the passband width of a low-pass filter affects its ISR results. A too wide passband or a too narrow passband can result in a decline in performance.
Though a low-pass filter can achieve ISR (please refer to Sec. 3), its performance can never surpass a well-trained neural network. The outcome of a low-pass filter varies with respect to the passband width, as depicted in Fig. 5. However, information outside the passband will be completely wiped out, causing an observable detail loss in high-frequency components. On the contrary, the non-linear part of neural networks is able to inject information in high-frequency domain based on learned or structural priors. Moreover, it can compensate the grid-like distortions brought by the windowed low-pass filter. Together with the linear part, neural networks function as the superset of low-pass filter, retaining both high and low frequency information.
We compute the non-linear response and its frequency spectrum of the neural network using the proposed HyRA paradigm. In the toy example presented in Fig. 4, it can be noticed that the response of the non-linear component exhibits sharper edges. Compared with the frequency spectrum of the ISR results, the central bright spot in the response of $G(I)$ spreads to a larger range, indicating that more power is distributed into the high-frequency domain. Almost all the components of the high-frequency part in the final ISR result are contributed by the non-linear component.
As mentioned in Section 4.1.1, the non-linear component also plays a crucial role in compensating for the distortion introduced by $H(I)$. Examining the response of $G(I)$, we note that it also exhibits grid-like distortions, matching those in the response of $H(I)$. This allows for the cancellation of the grid-like distortions, achieving the final goal of ISR. As shown in Fig. 4, upon observing the frequency spectrum, bright spots corresponding to the amplitude spectrum of $H(I)$ exist in all four corners of the amplitude spectrum of $G(I)$. However, the phase spectrum of $G(I)$ is in compensation of the phase spectrum of $H(I)$, indicating that the grid-like distortion is ‘erased’ here.
In summary, the non-linear component $G(I)$ serves to inject high-frequency details learned during training to compensate for the loss of high frequencies introduced by the low-pass filter. Simultaneously, it addresses distortions arising from the imperfect performance of the low-pass filter.
4.2 Frequency Spectrum Distribution Similarity (FSDS)
In this section, we introduce the FSDS metric to quantitatively describe the so called the ‘injected high frequencies’ as discussed in Section 4.1.2.
4.2.1 Motivation and Method
Figure 6: X-FFT-$\Sigma$ denotes the integrated frequency spectrum, the integration path is from origin to infinty in every quadrant. Columns 1 and 2 in the figure respectively show that the differences in the results of different ISR methods can be reflected in the frequency spectrum. Column 3 presents the integral of the spectrum from low to high frequencies in a contour plot. The distribution of contours visually represents the distinct distribution of different frequency components.
Since we need to measure the components of injected high frequencies, we must delve into the issue from a frequency spectrum perspective. However, commonly used metrics such as PSNR, SSIM [26], and LPIPS [36] do not measure the quality of an image from a spectral perspective.
Additionally, we’ve noted that the frequency domain distribution in the ISR field can significantly impact downstream applications [35, 28]. Consequently, we propose that evaluating the ISR effectiveness of a network requires a thorough assessment of its performance in the frequency spectrum. This involves examining the similarity in frequency spectrum between the low-resolution image and the high-resolution image. The Frequency Spectrum Distribution Similarity (FSDS) metric integrates the power distribution maps of the spectrum for both images. The difference is then calculated to generate an error map, and the total sum of its absolute values is computed.
For an image $I^{\mathrm{HR}}_{x,y}$, to minimize the impact of the data input range on the results, we normalize the input data and then perform a two-dimensional Fourier transform to obtain $I^{\mathrm{HR}}_{j\omega_{1},j\omega_{2}}$, which can be mathematically described as:
$I^{\mathrm{HR}}_{j\omega_{1},j\omega_{2}}=\mathcal{F}\left[\frac{I^{\mathrm{HR}}-E(I^{\mathrm{HR}})}{\sigma(I^{\mathrm{HR}})}\right],$ (7)
where $E(I^{\mathrm{HR}})$ and $\sigma(I^{\mathrm{HR}})$ are the mean value and variance of $I^{\mathrm{HR}}$ respectively. Similarly, we perform a Fourier transform on the ISR image to obtain $I^{\mathrm{SR}}_{j\omega_{1},j\omega_{2}}$. It is worth noting that unlike other metrics, such as PSNR and SSIM[26], which do not incorporate normalization, FSDS is specifically designed to accentuate numerical variations due to its emphasis on numerical changes rather than absolute numerical values. Then, the complex integration of the two spectrum is performed, providing the power distribution map $D^{\mathrm{HR}}$, which is defined as:
$D^{\mathrm{HR}}=\iint_{(\omega_{1},\omega_{2})\in\mathbb{R}^{2}}I^{\mathrm{HR}}\mathrm{d}\omega_{1}\mathrm{d}\omega_{2}.$ (8)
Similarly, we can obtain $D^{SR}$. Subsequently, the difference between $D^{HR}$ and $D^{SR}$ is calculated, providing a difference map $D^{\mathrm{diff}}$ of their power distribution:
$D^{\mathrm{diff}}=D^{\mathrm{HR}}-D^{\mathrm{SR}}.$ (9)
Finally, we define the frequency spectrum distribution similarity (FSDS) as:
$\operatorname{FSDS}=-10\log_{10}\frac{\iint_{(\omega_{1},\omega_{2})\in\mathbb{R}^{2}}|D^{\mathrm{diff}}|^{2}\mathrm{d}\omega_{1}\mathrm{d}\omega_{2}}{\iint_{(\omega_{1},\omega_{2})\in\mathbb{R}^{2}}|D^{\mathrm{HR}}|^{2}\mathrm{d}\omega_{1}\mathrm{d}\omega_{2}},$ (10)
where ||$\cdot$|| ⋅ | represents taking the magnitude of a complex number. Considering a more concise description of a larger dynamic range, logarithm is taken. A larger FSDS value indicates that the two images are closer, thereby suggesting better ISR results.
4.2.2 The merits of FSDS
Figure 7: A comparison of SSIM and FSDS in JPEG compression and steganography. As can be seen, SSIM fails to reflect distortion brought by steganography, while FSDS captures both cases of distortion.
Previous image evaluation metrics, such as PSNR, SSIM [26], have focused on statistical or structural features of images, but no work has evaluated images from the perspective of their frequency spectrum. The spectrum is the concentrated expression of components with different changing rates in a signal or image. It is crucial for capturing details, eliminating noise, and comprehensively understanding image features. In image processing, spectrum analysis provides a more accurate evaluation, particularly playing a key role in applications sensitive to details. Due to the nature of Fourier transformation, which involves every pixel of the image in the computation, it encompasses not only information such as signal-to-noise ratio and structural similarity but also the overall similarity of the entire image. Therefore, evaluating image quality from the perspective of the spectrum is highly reasonable and necessary. Our FSDS metric can reflect distribution differences by employing a paradigm of integrating first in the frequency spectrum and then comparing. In other words, FSDS not only reflects the signal-to-noise ratio captured by the PSNR metric and the structural similarity indicated by the SSIM metric, but also captures features that these two metrics cannot represent. In the next paragraph, we will use two toy examples to demonstrate the rationale and advantages of FSDS.
From Fig. 6, it can be observed that images obtained by different ISR methods have different proportions of high-frequency components (the center of the spectrum figure represents low frequency, while higher frequencies extend outward). After integration, this is reflected in the varying widths of the dark cross-shaped patterns in the center. A narrower width indicates a higher proportion of low-frequency components in the spectrum, and vice versa. Existing methods may not effectively capture the loss of high-frequency components with low power in the frequency spectrum. Performing information steganography in the frequency spectrum can effectively highlight this aspect. As shown in Fig. 7, we embed some content in the frequency spectrum of the image. Such steganography causes our FSDS metric to drop to 26.37dB while the SSIM metric remains in a high level of 0.995. we can observe that after applying specific steganography to the spectrum of an image, the image exhibits some blurring and oscillation. Such oscillations are actually the Gibbs phenomenon, a typical oscillation phenomenon caused by the loss of high-frequency information. Meanwhile, when we apply JEPG compression to the image [^5] In this example, the compression quality is set to 10., when FSDS drops to 26.39dB, SSIM together drops to 0.842. This toy example demonstrates that there indeed exists some feature SSIM cannot reflect while that can be reflected by FSDS.
In summary, previous methods may not effectively reflect the situation in the image frequency spectrum, while our proposed FSDS metric can sensitively detect distortions in the frequency spectrum.
5 Experiments
Figure 8: The ISR performance using a low-pass filter shows variations with the cutoff frequency $\omega$. This figure illustrates the results obtained from the $\times$2 ISR task conducted on the DIV2K dataset. To enhance the clarity of the visualization, the curve has been smoothed using a moving average with a window length of 10.
Figure 9: Comparison of impulse responses and the sinc function for several mainstream backbone networks and their derivatives. The impulse response of the bicubic interpolation result is presented as a reference.
Due to the page limitation, we can only present three of the most crucial experiments in this section, namely: 1) the relationship between the low-pass filter passband width and ISR performance; 2) various network impulse responses; 3) a comparison of FSDS metrics with PSNR, SSIM and LPIPS [36] metrics on the DIV2K dataset. For more experiments, please refer to Appx. C.
5.1 Experiment on Low-pass Filtering Super-resolution Performance
In Section 4.1.2, we mention that a vanilla low-pass filter can achieve ISR, we now present an experiment on the relationship between the low-pass filter passband width and ISR performance. As shown in Fig. 8, we utilized various low-pass filters to perform $\times$2 ISR on the validation set from of DIV2K dataset. Subsequently, we evaluated the ISR results using the PSNR and SSIM metrics. When $\omega=48$, PSNR reaches its maximum value of $31.40$. When $\omega=45.8$, SSIM reaches its maximum value of $0.87$. We assert that, in terms of neural network performance, for $\times$2 ISR, the PSNR should not fall below $31.40$, and the SSIM should not be lower than $0.87$. Otherwise, it can be considered that the neural network may not effectively capture both low-frequency and high-frequency information.
5.2 Experiment on Impulse Response
We select several mainstream backbones and their derivatives commonly used for the ISR task [13, 5, 20, 27, 38, 7, 12, 10, 37] and conduct impulse response tests. The experimental results are compared with the sinc function and depicted in Fig. 9. The input image is an $11\times 1$ 11 image where only the pixel at position $(5,5)$ is white (the values for all three channels at this position are 255, with indices starting from 0), and the rest of the image is black (with values of 0). According to Tab. 3, it can be observed that as the ISR factor increases, the central peak of the output sinc function becomes wider and more pronounced. To balance visual saliency and the maximum ISR factor achievable by certain networks, we opted for a 4x ISR factor. Observing the experimental results, we can notice that regardless of the neural network structure used for ISR, whether it’s a CNN or a transformer, the impulse response exhibits some degree of similarity to the two-dimensional sinc function. This similarity is particularly pronounced in networks like RDN [38] and RCAN [37]. Despite some distortion in comparison to the sinc function, EDSR [13], EQSR [25], and their derivatives still exhibit significant features of the sinc function, including the central bright spot and elongated bright patches in the cardinal directions. From Tab. 1, we observe that networks exhibiting superior performance tend to generate impulse responses that closely resemble the sinc function. This observation suggests that preserving low-frequency information more effectively can also enhance performance. However, few previous works has focus on low-frequency, giving us a new idea for furture ISR networks.
5.3 Experiment on FSDS Metric
Method PSNR SSIM LPIPS FSDS (Ours)
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$
EDSR[13] 34.6312 30.9514 28.8716 - - 0.93712 0.87413 0.81616 - - 0.04212 0.10114 0.15516 - - 39.2115 34.1510 31.387 - -
EDSR-LIIF[13] 34.5514 30.9215 28.9815 26.764 23.754 0.93715 0.87414 0.81914 0.7414 0.6334 0.04315 0.10013 0.15313 0.2434 0.4282 39.3713 34.536 31.329 28.454 23.153
EDSR-OPESR[13] 34.3416 30.9612 29.0412 - - 0.93616 0.87511 0.82013 - - 0.04313 0.10011 0.15314 - - 39.806 34.635 31.2911 - -
EDSR-SRNO[13] 34.729 31.0510 29.1310 26.903 23.873 0.9399 0.87610 0.82211 0.7463 0.6383 0.0418 0.09810 0.14910 0.2413 0.4374 39.5311 34.537 31.456 28.463 22.784
EDSR-LTE[13] 34.6113 30.9711 29.0314 - - 0.93714 0.87412 0.82012 - - 0.04314 0.10012 0.15212 - - 39.2914 34.338 31.3010 - -
RDN[38] 34.6910 30.5816 29.1211 - - 0.93810 0.86716 0.82310 - - 0.0419 0.10616 0.15011 - - 40.023 32.9516 31.654 - -
RDN-LIIF[38] 34.868 31.218 29.269 26.992 23.932 0.9398 0.8798 0.8269 0.7492 0.6392 0.04110 0.0968 0.1478 0.2311 0.4061 39.699 34.833 31.833 28.891 23.781
RDN-OPESR[38] 34.5215 31.199 29.288 - - 0.93811 0.8799 0.8268 - - 0.04211 0.0967 0.1489 - - 40.192 34.962 31.485 - -
RDN-LTE[38] 34.917 31.267 29.317 27.051 23.991 0.9397 0.8797 0.8277 0.7501 0.6411 0.0417 0.0956 0.1446 0.2332 0.4313 39.825 34.754 31.842 28.652 23.352
SwinIR-classical[12] 35.345 31.645 29.634 - - 0.9435 0.8855 0.8355 - - 0.0385 0.0924 0.1405 - - 40.371 35.131 32.371 - -
ITSRN[32] 32.6717 30.4917 28.7317 26.645 23.725 0.92217 0.86617 0.81317 0.7365 0.6305 0.05217 0.11317 0.16717 0.2715 0.4695 31.2518 26.1818 25.8818 25.625 21.575
HAT-S[4] 35.462 31.723 29.723 - - 0.9442 0.8873 0.8373 - - 0.0382 0.0923 0.1393 - - 39.787 33.8013 31.0614 - -
HAT[4] 35.462 31.772 29.752 - - 0.9442 0.8872 0.8372 - - 0.0382 0.0902 0.1382 - - 39.787 33.9111 31.2012 - -
HDSRNet[22] 34.6411 30.9513 29.0413 - - 0.93713 0.87315 0.81915 - - 0.04316 0.10315 0.15415 - - 39.4612 34.209 31.338 - -
GRLBase[11] 35.661 31.931 29.911 - - 0.9451 0.8891 0.8411 - - 0.0371 0.0891 0.1351 - - 39.994 33.8412 31.1213 - -
GRLSmall[11] 35.394 31.654 29.635 - - 0.9434 0.8864 0.8354 - - 0.0384 0.0925 0.1404 - - 39.5410 33.6814 31.0315 - -
GRLTiny[11] 35.176 31.416 29.406 - - 0.9426 0.8826 0.8306 - - 0.0396 0.0969 0.1467 - - 39.2016 33.2015 30.5616 - -
Bicubic 31.0418 28.2518 26.6918 24.876 22.346 0.89318 0.81318 0.75218 0.6756 0.5876 0.09618 0.19118 0.29118 0.4396 0.6136 32.7917 28.9017 26.5717 23.456 18.746
Table 1: Comparison of PSNR, SSIM [26], LPIPS [36] and FSDS metrics for different methods on the DIV2K dataset [1]. Items with the highest and the second-highest mean values are highlighted in red and blue, respectively. The gray superscripts are the order of each method.
We conducted tests on the validation set of the DIV2K dataset [1] using several methods [13, 5, 20, 27, 10, 38, 12, 32], depicted in Tab. 1. The evaluation metrics include PSNR, SSIM [26], LPIPS [36], and our FSDS. All tests are performed using code and weights available in open-source official repositories. For all methods, we conduct experiment for $\times$2 to $\times$4. For methods that support arbitrary-scale ISR, we test for $\times$6 and $\times$12 as well. In the $\times$2 to $\times$4 range, GRLBase[11] consistently achieves the best performance across PSNR, SSIM and LPIPS metrics, and FSDS shows that SwinIR [12] achieves the best performance. For $\times$6 and $\times$12, RDN-LTE [10] exhibits the best PSNR and SSIM metrics, while RDN-LIIF performs best on LPIPS and the FSDS metric.
From Section 4.2, we claim that previous metrics are not sensitive to high-frequency information, while FSDS does. This can be proven by Tab. 1. In the case of slight high-frequency loss, such as on scales $\times$ 2 to $\times$ 4, FSDS responds differently compared to previous metrics. In cases that suffer from severe high-frequency loss, such as on $\times$ 6 and $\times$ 12 scales, FSDS shows consistency with previous metrics. This is because when high-frequency loss is slight, previous metrics fail to reflect such high-frequency loss and while the loss becomes more severe, they start to capture such loss. This observation shows the necessity of applying the FSDS metrics to assess image quality objectively.
5.4 Some Exceptions to Impulse Responses
Figure 10: The impulse response of SwinIR-Real [12] and ESRGAN [24] is not an obvious sinc function.
We observe that not all impulse response of networks is ‘sinc’ function, as shown in Fig. 10. SwinIR-Real [12] and ESRGAN [24] are trained using adversarial loss, while methods in Fig. 9 uses loss like $\ell_{1}$ or $\ell_{2}$ loss. Therefore, we believe the ‘sinc’ impulse response is related to the loss function.
6 Conclusion
In this paper, we report an intriguing observation. i.e., the sinc phenomenon, which reveals that the impulse response of ISR networks act as low-pass filters. Building on this observation, we introduce a novel approach called Hybrid Response Analysis (HyRA) to explore the hidden behavior of ISR networks. HyRA treats a neural network as a combination of a linear system and a non-linear system with a zero impulse response. The linear system functions as a low-pass filter, while the non-linear system utilizes prior knowledge to inject high-frequency details. To assess the neural network’s information recovery across the frequency spectrum, we propose the Frequency Spectrum Distribution Similarity (FSDS) metric. FSDS uncovers properties overlooked by previous metrics, and experiments validate the rationality and necessity of it.
Acknowledgements
We appreciate anonymous reviewers for their previous suggestions to help this paper better. Moreover, we would like to express our sincere gratitude to Ruijie Zhu (rzhu48@ucsc.edu) for his generous support in GPUs. Without his support, it is hard for us to do experiments using full-scale DIV2K dataset. This work is supported by NSFC (12271083).
Impact Statement
This paper presents work whose goal is to advance the interpretability of neural networks in Image super-resolution. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Notation Conventions
Symbols
$j$ Imaginary number unit
* Convolution operator
$I_{x,y}^{\mathrm{comment}}$ 2-D signal with variant $x,y$
$I_{j\omega_{1},j\omega_{2}}^{\mathrm{comment}}$ Fourier transform of $I_{x,y}$
$x(t)$ 1-D signal with variant $t$
$X(j\omega)$ Fourier transform of $x(t)$, $j\omega$ is a notation, $\omega$ is the variant
$x[n]$ Discrete signal with index $n$
$X[k]$ DFT of $x[n]$
$\mathcal{F}[x(t)]$ Fourier transform operator, $X(j\omega)=\mathcal{F}[x(t)]$
$\mathcal{F}^{-1}[X(j\omega)]$ Inverse Fourier transform, $x(t)=\mathcal{F}^{-1}[X(j\omega)]$
Signals
$\delta(t)$ Dirac $\delta$ function
$sinc_{\omega}(t)$ The sinc funtion with parameter $\omega$, $sinc_{\omega}(t)=\frac{sin(\omega t)}{\pi t}$. The sinc function is the time-domain waveform of an ideal low-pass filter.
$sinc^{\omega}_{x,y}$ 2-D sinc function with parameter $\omega$, $sinc^{\omega}_{x,y}=\frac{sin(\omega x)}{\pi x}\cdot\frac{sin(\omega y)}{\pi y}$
$s_{\Delta T}(t)$ 1-D sample signal with a sample interval of $\Delta T$, $s_{\Delta T}(t)=\sum\limits_{n=-\infty}^{\infty}\delta(t-nT)$
Table 2: Notation Conventions
Appendix B Signal Processing Theories
We briefly introduce some related concepts and methods used in this paper in this section.
B.1 System and Response
The word ‘system’ has many meanings and interpretations. This paper views a system as a process in which input signals are transformed by the system or cause the system to respond in some way, resulting in other signals as output [18]. Systems can be divided into linear systems and nonlinear systems according to their mathematical properties. A linear system refers to a system with such a property: the response of the system to the input $x_{1}(t)$, $x_{2}(t)$ is $y_{1}(t)$, $y_{2}(t)$ respectively, then when the input is $x_{1}(t)+x_{2}(t)$, the response of the system is $y_{1}(t)+y_{2}(t)$.
Systems can also be divided into time-variant ones and time-invariant ones according to their temporal properties. A time-invariant system refers to that the properties of the system do not change with time, that is, the system has the same impulse response at any time. It satisfies such a relationship: when the input is $x(t)$, the output is $y(t)$, and when the input is $x(t-t_{0})$, the output is $y(t-t_{0})$.
A system with both linear and time-invariant properties is a linear time-invariant (LTI) system. For an LTI system, we can use ‘impulse response’ to uniquely describe it: systems with the same impulse response are the same system, vice versa. The impulse response $h(t)$ is defined as the output of the system when the input signal is $\delta(t)$ (Dirac delta function). The response of a linear system to an arbitrary input signal can be computed through the convolution operation of its impulse response and the input signal, namely:
[…]
In the equation, * is the convolution operator, $y(t)$ is the system output and $x(t)$ is the input signal. When we apply Fourier transform to the impulse response $h(t)$, then we can obtain the transfer function $H(j\omega)$ of the system. The transfer function describes the frequency domain waveform of the impulse response. According to the convolution theorem, the response of a linear system can also be obtained by multiplying the Fourier transform of the input signal by the transfer function of the system and then performing the inverse Fourier transform. In summary, given the impulse response of an LTI system, we can calculate the system’s response to any output.
B.2 Signal Sampling and Recovery
Figure 11: Time-domain to frequency-domain waveform variation of the continuous signal sampling process. The sampling function $s_{\delta T}$ is an impulse train sequence with an interval of $T$, and $S_{\delta T}(t)$ is its frequency domain waveform, which is also an impulse train sequence. Sampling a signal causes duplication in the frequency domain.
Figure 12: Time-domain to frequency-domain waveform variation in the process of sampling signal recovery. $h_{LP}(t)$ is the time-domain impulse response of a low-pass filter, and $H_{LP}(j\omega)$ is its frequency-domain waveform.
Signal sampling and sample recovery are very common operations, and in this section, we will briefly analyze this process from the perspective of both the time-domain and frequency-domain. The upper part of Fig. 11 shows the time domain waveform variation of signal sampling process, and the lower part shows the frequency domain waveform variation of signal sampling process. To sample a continuous signal, the sampling process can be regarded as the multiplication of the original signal $x(t)$and an impulse train signal $s_{\Delta T}(t)$. It can be described as:
$\left\{\begin{array}[]{c}x^{\prime}(t)=x(t)\cdot s_{\Delta T}(t),\\ s_{\Delta T}(t)=\sum\limits_{n=-\infty}^{\infty}\delta(t-nT),\end{array}\right.$ (12)
where $T$ denotes the sampling interval, $x^{\prime}(t)$ denotes the sampled signal. According to the convolution theorem, the frequency domain change of the sampling process can be described in the following way:
$X^{\prime}(j\omega)$ $=X(j\omega)*S_{\delta T}(j\omega)$ (13)
$=\sum\limits_{k=-\infty}^{\infty}X[j(\omega-k\frac{2\pi}{T})].$
That is, the sampling process is reflected in the frequency spectrum as a periodic extension of the frequency spectrum of the original signal $x(t)$.
Fig. 12 shows the time domain and frequency domain waveform variation during the recovery process. For sampling recovery, in order to restore the sampled signal $x^{\prime}(t)$ to the original signal $x(t)$, from the perspective of frequency domain, a low-pass filter is all we needed, that is, convolving the sampled signal with a low-pass filter $h_{LP}(t)$. This process can be expressed as:
$x(t)$ $=x^{\prime}(t)*h_{LP}(t)$ (14)
$=x^{\prime}(t)*sinc_{\omega}(t),$
where in the equation, $h_{LP}(t)=sinc_{\omega_{0}}(t)=\frac{sin(\omega_{0}t)}{\pi t}$ is the time domain response of the ideal low-pass filter, and its frequency-domain waveform $H_{LP}(j\omega)$ is a rectangular window.
B.3 Spectrum Aliasing
Figure 13: The illustration of spectrum aliasing. On the left, there is no aliasing as the sampling rate is sufficiently high. On the right, aliasing occurs due to an insufficient sampling rate.
Spectrum aliasing is a manifestation of information loss. Fig. 13 depicts the time-domain and frequency-domain scenarios of no frequency overlapping and frequency overlapping, respectively. When the sampling rate is lower than the Nyquist sampling rate [^6] The minimum sampling rate that can completely restore the origin of the sampled from sampled signal, which is twice the highest frequency of the original signal [16]. When the sampling rate is below the Nyquist sampling rate, the approach mentioned in Section B.2 cannot completely restore the original signal $x(t)$. From Tab. 3, we can see that for the sample signal $s_{\Delta T}(t)$, the larger $T$ is, the sparser its time domain impulse train gets, while in the frequency spectrum the impulse trains gets denser. When the impulse trains in the frequency domain become sufficiently dense, and the spectrum of the original signal is periodically extended, overlapping occurs, preventing the complete recovery of the original signal. In ISR tasks, spectrum aliasing is manifested when restoring a low-resolution image to a high-resolution image, resulting in the loss of high-frequency information such as details and textures.
Appendix C Extra Experiments
C.1 Linear and Non-linear Responses
Figure 14: Linear and non-linear responses and their corresponding frequency spectrum of various ISR methods.
In Fig. 14, we present the linear and nonlinear responses of various ISR networks along with their corresponding spectrums. From the figure, it is evident that different networks exhibit varying filtering effects in their linear components. EDSR demonstrates a pronounced removal of high-frequency components, and compared to other methods, it exhibits the smallest area of brightness diffusion around the central bright spot in its spectrum. From the nonlinear responses, it can be observed that the nonlinear components of the networks are all involved in supplementing high-frequency information and correcting distortions.
C.2 Space Invariance
Figure 15: Spatial invariance experiment conducted on SwinIR [12]. When we feed the SwinIR network with impulses at various positions, the ISR results demonstrate that the RDN exhibits spatial invariance.
We conducted spatial invariance testing on RDN [38] (For the concept of spatial invariance, please refer to Section 4.1). The input data consists of an image where only one pixel is white (the pixel value is 1), and all other pixels are black (the pixel value is 0). By shifting the position of this white pixel, we obtain $I(x-\Delta x,y-\Delta y)$. This shifted input $I(x-\Delta x,y-\Delta y)$ is then fed into the neural network, and we obtain its shifted impulse response, as illustrated in Fig. 15. Observing the experimental results, we find that the responses to different $I(x-\Delta x,y-\Delta y)$ are consistent, with the only difference being their position. This demonstrates that, for ISR networks, the linear component in HyRA exhibits spatial invariance.
C.3 Exploration of the Positional Origin of Sinc-like Patterns
Figure 16: Visualization of feature maps in different EDSR [13] network layers. The sinc-like pattern start to take shapes after the sub-pixel convolution, before the last convolution layer.
As shown in Fig. 16, we visualize the output features of different components in the EDSR [13] network for analysis. We observe that the approximate shape of the sinc function begins to take form after the Upsampler module, and after a convolution, it essentially forms the shape of a sinc function. Interestingly, in the EDSR network, the Upsample module uses sub-pixel convolution (convolution + pixel shuffle) for upsampling without any interpolation. This indicates that the low-pass filter present in the network is learned by the network itself and not introduced by interpolation kernels.
Appendix D The Fourier Transform Pairs Involved in This Paper
Symbol/Name Section(s) Time domain Frequency Domain
$s_{\Delta T}(t)$, $\Delta T$ is the samping interval Section B.2, Section B.3 $\sum\limits_{n=-\infty}^{\infty}\delta(t-nT)$ $\frac{2\pi}{T}\sum\limits_{k=-\infty}^{\infty}\delta(\omega-k\frac{2\pi}{T})$
Ideal Low-pass filter Section 4.1.1 $x(t)=sinc_{\omega_{0}}(t)=\frac{sin(\omega_{0}t)}{\pi t}$, $\omega_{0}$ is called the cut-off frequency $X(j\omega)=\begin{cases}1,&|\omega|<\omega_{0}\\ 0,&|\omega|>\omega_{0}\end{cases}$
$\delta(t)$ Section 4.1.1 $\lim\limits_{\tau\rightarrow 0}\int_{-\tau}^{+\tau}\delta(t)=1$ 1
Table 3: Fourier transform pairs
Appendix E The Windowing Operation
Figure 17: Various window functions.
The time-domain waveform of an ideal low-pass filter is a sinc function. The sinc function is defined over [,][$\infty$,$\infty$][ - ∞ , ∞ ], and the number of zero crossings is countable. This implies that in reality, an ideal low-pass filter does not exist. In discrete-time signal processing, truncating a designed filter using a window function is common. There are many window functions, such as the rectangular window, Hanning window, Blackman window, and so on. Fig. 17 illustrates some commonly used window functions. Observing the experimental results and analyzing the relationship between the peak values of the main lobe and the first side lobe, we find that the impulse response of the neural network seems to undergo windowing. However, different networks appear to adopt different window functions.
Appendix F Frequency Spectrum Period Extension Caused by Zero Padding
Figure 18: Performing zero-padding on an image to reach the target size will result in periodic extension in the frequency spectrum obtained through its Discrete Fourier Transform.
When considering integer factor ISR, our approach to computing the linear component response is as follows: first, upsample the low-resolution image to the high-resolution image through zero-padding, and then convolve it with the impulse response to obtain the response. During the zero-padding process, it leads to period extension in the frequency spectrum. For a signal $x[n]$ of length $N$ undergoing DFT to obtain $X[k]$, we have:
$X[k]=\sum\limits_{n=0}^{N-1}x[n]e^{-j\frac{2\pi}{N}kn}.$ (15)
Then, zero-padding is applied to $x[n]$, producing in a new signal $x_{2}[n]$ of length $3N$:
$x_{2}[n]=\left\{\begin{array}[]{ll}x[\frac{n}{3}],&n=0,3,\cdots,3N-3\\ 0,&otherwise.\end{array}\right.$ (16)
Perform DFT to $x_{2}[n]$ to obtain $X^{\prime}[k]$, then we have:
$X^{\prime}[k]&=\sum\limits_{n=0}^{3N-1}x_{2}[n]e^{-j\frac{2\pi}{3N}k\cdot n}\\ &=\sum\limits_{n=0}^{N-1}x[n]e^{-j\frac{2\pi}{3N}k\cdot 3n}\\ &=\sum\limits_{n=0}^{N-1}x[n]e^{-j\frac{2\pi}{N}kn}\\ \end{aligned}.$ (17)
When $k<N$, there exists:
$e^{-j\frac{2\pi}{N}kn}=e^{-j\frac{2\pi}{N}(k+N)n}=e^{-j\frac{2\pi}{N}(k+2N)n}=\cdots.$ (18)
Therefore,
$X^{\prime}[k]=\left\{\begin{array}[]{ll}X[k]&0\leq k<N\\ X[k\ mod\ N]&N\leq k<3N-1\end{array}\right.$ (19)
Thus, zero-padding causes period extension in the frequency spectrum. Ideally, the extended spectrum would be filtered out by the low-pass filter used in the ISR process. However, due to the limited filtering capability of the potential filters within the neural network, the stopband attenuation is low, and the extended spectrum cannot be completely filtered out.
Appendix G Comparison Between FSDS and $\ell_{1}$, $\ell_{2}$ norms
We compare our proposed FSDS metric with $\ell_{1}$ norm, and $\ell_{2}$ norm on both frequency domain and image domain as depicted in Tab. 4 and Tab. 5. From these two figures, we can observe that $\ell_{1}$ norm and $\ell_{2}$ norm produces similar ranking orders when they are calculated on the same domain, indicating that $\ell_{1}$ norm is equivalent to $\ell_{2}$ norm when assessing image quality. However, the ranking orders produced by FSDS is distinctive to that of $\ell_{1}$ and $\ell_{2}$. This means our FSDS metric reflects image quality in a unique way.
Method FSDS (Ours) $\ell_{1}$ Norm in Frequency Domain $\ell_{1}$ Norm
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$
EDSR[13] 39.21015 34.14810 31.3807 - - 35.19012 51.67612 62.47816 - - 0.01210 0.01913 0.02316 - -
EDSR-LIIF[13] 39.37113 34.5356 31.3219 28.4464 23.1473 35.36914 51.88614 61.99615 73.9634 88.2035 0.01315 0.01915 0.02315 0.0294 0.0424
EDSR-OPESR[13] 39.7986 34.6305 31.28611 - - 36.79616 51.94815 61.71114 - - 0.01316 0.01811 0.02312 - -
EDSR-SRNO[13] 39.53311 34.5277 31.4486 28.4583 22.7784 34.68810 51.11310 61.14011 73.1363 87.7043 0.01211 0.01810 0.02310 0.0293 0.0413
EDSR-LTE[13] 39.29014 34.3288 31.30310 - - 35.11011 51.55611 61.70612 - - 0.01314 0.01914 0.02314 - -
RDN[38] 40.0223 32.94616 31.6464 - - 34.5959 53.43116 61.06910 - - 0.01312 0.01916 0.02311 - -
RDN-LIIF[38] 39.6909 34.8313 31.8323 28.8941 23.7781 34.2248 50.4348 60.5639 72.7532 87.4882 0.0128 0.0189 0.0229 0.0292 0.0412
RDN-OPESR[38] 40.1882 34.9592 31.4755 - - 36.12615 50.7929 60.3888 - - 0.01313 0.0187 0.0227 - -
RDN-LTE[38] 39.8255 34.7494 31.8372 28.6542 23.3462 33.9977 50.1067 60.2297 72.3011 87.0901 0.0127 0.0188 0.0228 0.0291 0.0411
SwinIR-classical[12] 40.3721 35.1251 32.3701 - - 32.7505 48.3804 58.5794 - - 0.0125 0.0175 0.0215 - -
ITSRN[32] 31.25418 26.17818 25.87618 25.6195 21.5665 41.92817 53.45917 62.92717 74.2605 88.1604 0.01517 0.02017 0.02417 0.0305 0.0425
Bicubic 32.79017 28.89617 26.56817 23.4506 18.7366 50.57118 64.94918 73.41218 82.5676 93.0126 0.01818 0.02418 0.02918 0.0366 0.0506
HAT-S[4] 39.7847 33.79913 31.05714 - - 32.3872 48.0473 58.2323 - - 0.0112 0.0173 0.0213 - -
HAT[4] 39.7847 33.91311 31.19612 - - 32.3872 47.8152 58.0522 - - 0.0112 0.0172 0.0212 - -
HDSRNet[22] 39.45812 34.2039 31.3258 - - 35.24513 51.69013 61.70913 - - 0.0129 0.01912 0.02313 - -
GRLBase[11] 39.9904 33.84012 31.12113 - - 31.8071 47.2631 57.2961 - - 0.0111 0.0171 0.0211 - -
GRLSmall[11] 39.54410 33.67914 31.02715 - - 32.7144 48.5935 58.8055 - - 0.0124 0.0174 0.0214 - -
GRLTiny[11] 39.20516 33.19715 30.55616 - - 33.3946 49.5506 59.8356 - - 0.0126 0.0186 0.0226 - -
Table 4: Comparison between our proposed FSDS metric and $\ell_{1}$ norm in both frequency domain and image domain. Items with the highest mean values are highlighted in red and secondary mean values in blue. The gray superscripts denote the ranking order.
Method FSDS (Ours) $\ell_{2}$ Norm in Frequency Domain $\ell_{2}$ Norm
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$$\times 2$$\times 3$$\times 4$$\times 6$$\times 12$
EDSR[13] 39.21015 34.14810 31.3807 - - 4966.90513 11163.14815 17206.00916 - - 4966.90613 11163.14915 17206.01016 - -
EDSR-LIIF[13] 39.37113 34.5356 31.3219 28.4464 23.1473 4967.72714 11163.11214 16947.23515 27007.7484 49280.8834 4967.72714 11163.11314 16947.23615 27007.7494 49280.8864
EDSR-OPESR[13] 39.7986 34.6305 31.28611 - - 5247.22916 11117.69412 16818.59814 - - 5247.22916 11117.69512 16818.59914 - -
EDSR-SRNO[13] 39.53311 34.5277 31.4486 28.4583 22.7784 4794.9039 10845.39010 16453.70110 26236.2463 48097.2463 4794.9039 10845.39110 16453.70210 26236.2473 48097.2493
EDSR-LTE[13] 39.29014 34.3288 31.30310 - - 4907.85511 11032.25211 16781.62113 - - 4907.85511 11032.25211 16781.62213 - -
RDN[38] 40.0223 32.94616 31.6464 - - 4820.54010 12055.74516 16462.59411 - - 4820.54110 12055.74616 16462.59511 - -
RDN-LIIF[38] 39.6909 34.8313 31.8323 28.8941 23.7781 4627.8068 10477.3038 15993.7978 25720.7662 47650.6812 4627.8068 10477.3048 15993.7988 25720.7682 47650.6832
RDN-OPESR[38] 40.1882 34.9592 31.4755 - - 5061.53515 10609.1689 16034.9259 - - 5061.53515 10609.1699 16034.9269 - -
RDN-LTE[38] 39.8255 34.7494 31.8372 28.6542 23.3462 4576.6527 10349.2737 15808.8907 25384.5971 46955.7291 4576.6527 10349.2747 15808.8917 25384.5981 46955.7311
SwinIR-classical[12] 40.3721 35.1251 32.3701 - - 4222.2265 9617.2125 14804.5335 - - 4222.2265 9617.2125 14804.5345 - -
ITSRN[32] 31.25418 26.17818 25.87618 25.6195 21.5665 7483.58217 12244.89717 17878.39917 27664.0545 49417.0615 7483.58217 12244.89817 17878.40017 27664.0565 49417.0645
Bicubic 32.79017 28.89617 26.56817 23.4506 18.7366 10873.90018 19564.47318 26817.58018 38541.6756 64011.6986 10873.90118 19564.47418 26817.58118 38541.6776 64011.7026
HAT-S[4] 39.7847 33.79913 31.05714 - - 4107.3822 9423.4473 14498.5093 - - 4107.3832 9423.4483 14498.5103 - -
HAT[4] 39.7847 33.91311 31.19612 - - 4107.3822 9326.6132 14407.4572 - - 4107.3832 9326.6132 14407.4582 - -
HDSRNet[22] 39.45812 34.2039 31.3258 - - 4948.23712 11121.91213 16697.18512 - - 4948.23812 11121.91213 16697.18612 - -
GRLBase[11] 39.9904 33.84012 31.12113 - - 3937.0131 9073.9131 13959.0781 - - 3937.0131 9073.9131 13959.0791 - -
GRLSmall[11] 39.54410 33.67914 31.02715 - - 4170.8534 9585.3264 14742.7874 - - 4170.8534 9585.3274 14742.7884 - -
GRLTiny[11] 39.20516 33.19715 30.55616 - - 4384.3646 10079.0176 15488.3166 - - 4384.3646 10079.0186 15488.3176 - -
Table 5: Comparison between our proposed FSDS metric and $\ell_{2}$ norm in both frequency domain and image domain. Items with the highest mean values are highlighted in red and secondary mean values in blue. The gray superscripts denote the ranking order. To demonstrate Parseval’s theorem, we omit the mean operation when calculating $\ell_{2}$ norm.
Appendix H How G(I) is Trained and How G(I) and H(I) Vary Dependently
Figure 19: $N(I)$, $H(I)$ and $G(I)$ from different epoches during training.
To better explain how $G(I)$ is trained and how $G(I)$ and $H(I)$ vary dependently, we conduct a new experiment on observing their varying progress during training. We train a vanilla RDN [38] network from scratch and obtain the impulse response, $H(I)$, $G(I)$, and $N(I)$ from each epoch (see Fig. 19). As shown in the figure, in the second row, the $sinc$ phenomenon becomes clearer along with the training process, this indicates that the network is gradually learning the low-pass filter. The same conclusion can be further supported by observing the variation in $H(I)$ as shown in the third row. In this row, $H(I)$ illustrates the phenomenon of the grid-like distortion vanishing while low-frequency areas getting smoother. The fourth row depict the plot of $G(I)$. This row demonstrates that $G(I)$ is capturing more and more high-frequency information, such as edges. This is very significant especially when comparing the result from Epoch 1 and Epoch 50. Such increase in high-frequncy information can also be found in the frequency spectrum. During the entire training process, the network fails to recover the words on the wall (please see the magnified area, there are words on the wall next to the blue awning). Observing $G(I)$, we can find no sign of words as well. This indicates that the network treats it as low-frequency information and ignores it, pointing the way for future network improvements.
Appendix I Code Repositories
Abbreviate Title Publication Year Code Link
EDSR [13] Enhanced Deep Residual Networks for Single Image Super-Resolution CVPRW 2017 Github
LIIF[5] Learning Continuous Image Representation with Local Implicit Image Function CVPR 2021 Github
OPE-SR[20] OPE-SR: Orthogonal Position Encoding for Designing a Parameter-Free Upsampling Module in Arbitrary-Scale Image Super-Resolution CVPR 2023 Github
SRNO[27] Super-Resolution Neural Operator CVPR 2023 Github
LTE [10] Local Texture Estimator for Implicit Representation Function CVPR 2022 Github
RDN [38] Residual Dense Network for Image Super-Resolution CVPR 2018 Github
SwinIR [12] SwinIR: Image Restoration Using Swin Transformer ICCV 2021 Github
ITSRN [32] Implicit Transformer Network for Screen Content Image Continuous Super-Resolution NeurIPS 2021 Github
RCAN [37] Image Super-Resolution Using Very Deep Residual Channel Attention Networks ECCV 2018 Github
HAT [4] Activating More Pixels in Image Super-Resolution Transformer CVPR 2023 Github
HDSRNet [22] Heterogeneous Dynamic Convolutional Network in Image Super-Resolution Arxiv 2024 Github
GRL [11] Efficient and Explicit Modelling of Image Hierarchies for Image Restoration CVPR 2023 Github
Table 6: The papers and repository links used in this paper.